OpenStudy (bloomlocke367):

A multiple choice has 10 questions. Each question has four answer choices. a. What is the probability a student randomly guesses the answers and gets exactly six questions correct? b. Is getting exactly 10 questions correct the same probability as getting exactly zero correct? explain. c. Describe the steps needed to calculate the probability of getting at least six questions correct if the student randomly guesses.

OpenStudy (bloomlocke367):

@rational

OpenStudy (bloomlocke367):

for a, is it 1/4*6/10?

OpenStudy (bloomlocke367):

@rational

OpenStudy (bloomlocke367):

@phi @Callisto @asnaseer

OpenStudy (bloomlocke367):

Can you help @phi?

OpenStudy (phi):

this sounds like you should use binomial distribution https://en.wikipedia.org/wiki/Binomial_distribution#Probability_mass_function Have you studied that ?

OpenStudy (bloomlocke367):

yes, but I never understood.

OpenStudy (phi):

There are different levels of understanding. In this case, at least try to write down the formula. can you do that ?

OpenStudy (bloomlocke367):

\[P=\left(\begin{matrix}N \\ x\end{matrix}\right)p^x(1-p)^{(N-x)}\]

OpenStudy (bloomlocke367):

@phi

OpenStudy (bloomlocke367):

@rational can you please come help?

OpenStudy (rational):

Yes thats the formula for part a, first figure out the values of N, x, and p

OpenStudy (bloomlocke367):

what do they mean? I don't know what they stand for...

OpenStudy (rational):

N = total number of times an experiment was conducted (10 here) x = number of successes (6 here) p = success probability (1/4 here)

OpenStudy (bloomlocke367):

ohhh

OpenStudy (bloomlocke367):

\[P=\left(\begin{matrix}10 \\ 6\end{matrix}\right)(1/4)^6(1-(1/4))^{(10-6)}\]

OpenStudy (bloomlocke367):

right?

OpenStudy (rational):

Yes evaluate

OpenStudy (bloomlocke367):

\[P=\left(\begin{matrix}10 \\ 6\end{matrix}\right)(1/4096)(3/4)^4\]

OpenStudy (bloomlocke367):

\[p=\left(\begin{matrix}10 \\ 6\end{matrix}\right)(1/4096)(81/256)\]

OpenStudy (bloomlocke367):

\[P=\left(\begin{matrix}10 \\ 6\end{matrix}\right)81/1048576\]

OpenStudy (bloomlocke367):

right?

OpenStudy (rational):

Yes use calculator

OpenStudy (bloomlocke367):

to do what?

OpenStudy (bloomlocke367):

?...

OpenStudy (bloomlocke367):

what did I do wrong?

OpenStudy (rational):

nothing, your work looks good

OpenStudy (rational):

just evaluate \(\large \binom{10}{6}\)

OpenStudy (bloomlocke367):

how?

OpenStudy (rational):

it is called binomial coefficient : \[\large \binom{n}{r} = \dfrac{n!}{r!(n-r)!}\]

OpenStudy (rational):

you must be familiar with factorials right ?

OpenStudy (bloomlocke367):

yes.

OpenStudy (bloomlocke367):

okay.

OpenStudy (bloomlocke367):

what about b? like I know, but I dunno how to explain it.

OpenStudy (bloomlocke367):

@rational

OpenStudy (rational):

for part b, work both probabilities and see ?

OpenStudy (bloomlocke367):

well I know it's different because to get it wrong you have 3/4 but to get it right you have 1/4, but I don't know how to explain it.

OpenStudy (bloomlocke367):

nevermind. I got it.

OpenStudy (rational):

`b. Is getting exactly 10 questions correct the same probability as getting exactly zero correct? explain.` probability for getting exactly 10 questions correct \[P(X = 10) ~~~=~~~ \binom{10}{10} (1/4)^{10}(1-1/4)^{10-10}\] probability for getting exactly 0 questions correct \[P(X = 0) ~~~=~~~ \binom{10}{0} (1/4)^{0}(1-1/4)^{10-0}\]

OpenStudy (rational):

do they both evaluate to same number ?

OpenStudy (bloomlocke367):

no

OpenStudy (bloomlocke367):

thank you... I got that part... now for the hard one, if you're still willing to help, C.

OpenStudy (rational):

it is not hard as we don't need to do any arithmetic the question is just asking us to explain how you would work it, its not asking u to actually work it

OpenStudy (bloomlocke367):

I guess so.

OpenStudy (rational):

`c. Describe the steps needed to calculate the probability of getting at least six questions correct if the student randomly guesses.` so the student should get >= 6 questions correct yes ?

OpenStudy (rational):

that is he can get 6 or 7 or 8 or 9 or 10 questions correct so you simply find the probbaility for each like you did in part a and add up

OpenStudy (bloomlocke367):

oh, okay... thank you so much.

OpenStudy (bloomlocke367):

I really appreciate it.

OpenStudy (rational):

P(X >=6) := P(X=6) + P(X=7) + P(X=8) + P(X=9) + P(X=10)

OpenStudy (bloomlocke367):

yeah, I get it now.

OpenStudy (bloomlocke367):

thanks.

OpenStudy (rational):

yw